As a new electrical technician, you are designing a large solenoid to produce a uniform 0.150-T magnetic field near the center of the solenoid. You have enough wire for 3000 circular turns. This solenoid must be 58.0 cm long and 2.80 cm in diameter.

Part A

What current will you need to produce the necessary field?
Answer #1

B = 0.15T

N = 3000

L = 0.58m

Since B = uNI/L

=) I = BL/uN = 0.15×0.58/4.pai×10^-7 × 3000 = 23.06A

=) Current I = 23.06A